CRY+VIS Interface Meeting
Date: 2021/10/29 13:00-13:50
Participants: Takahashi, Ikeda, Fabian, Yasui, Hirata, Aso, Yokozawa, Ushiba, Akutsu, Sato
- IX: Health check. Repair of the BF actuator. Replacement of HPCD. TF measurement of the payload.
- EX: Health check. Investigation of the BF damper (V). Replacement of HPCD. TF measurement of the payload.
- Schedule: Replacement of the LVDT card and investigation on the WAB on 11/2.
- IY: LVDT driver for IP/F0. H4 geophone.
- Replacement of Flip cable: It will start from 11/5.
- Mirror cleanliness: Report by Yasui
- TM height setting: ITMY
- Oplev: Setup of the TM Oplev in IY after the height adjustment.
- ACC: Ready to install to EX. Measurement for IY is planed on 11/1.
- Commissioning in air: Health check. Check of the IM PD in IX.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Nov. 25-26 in IXV.
- Heater control: PR2 and SR2 are ready. PR3, IY, EY remain.
Date: 2021/10/29 14:00-14:50
Participants: Takahashi, Sato, Ushiba, Fabian, Terrence, Sugiyama, Aso, Hirata, Ishizaki, Tanaka
- Repaired the BF actuator in IX.
- Replaced the LVDT driver for IP/F0 in IY. Calibrated the LVDTs.
- Installed the H4 geophone in IY.
- Investigated the vertical BF damper. The LVDT driver card should be replaced.
- We raised SRM by 2.8 mm in accordance to the expected beam height: it is at the correct position now.
- We measured the height of the BS to be 2.85 mm below the average height of the four flanges (Tomaru-san's marking).
- We moved the GAS filter keystones with coil-magnet actuation: F0 and BF were close to zero and F1 was alledgely close to zero.
F1 LVDT is not working well (entry 18191), so we checked the amount of actuation we used the last time we measured the height, when the LVDT was working, and used the same value (entry 17338 from the 30th of June, 2021). This is fine because we had the same temperature (23.6 °C) and the fishing rod had not been used.
- Then we raised the BS and measured it to be 0.3 mm below the reference.
- In SRM LVDT pin assignment error:
- We triple-checked the IP LVDT coil resistance values in order to discard any obvious errors.
- Then we opened SR3 and SRM chambers and compared the wiring: they are indeed different; SR3 is the correct one.
- Nevertheless, the IP LVDTs are still working!
The calibration curves look normal. See klog 7087.
We connected SRM IP LVDT #3 to SR3 IP LVDT #3 driver, and viceversa, SR3 IP LVDT #3 to SRM IP LVDT #3 driver. Then we excited the suspensions slightly and inspected the behaviour of the six LVDTs in an Ndscope screen. Everything seemed very ordinary. We could not find any strange behaviour.
- I'm not an LVDT expert, but I guess the LVDTs still work becuase of its symmertry.
Question: I tend to think we should change the pin assignment, but the fact that they LVDTs are still working makes me doubt about changing them, what do you think?
- Unification of the magnet assignment in the suspensions.
- Type-A team: BF damper in EX (11/2). ACC for IY (11/1).
- Type-B team: SF in BF, Flip cables (11/5), Wind shield
(Date in Kamioka)
The week from 11/1
- Takahashi: 11/1-2
- Hirata: 11/2-5
The week from 11/8
- Takahashi: 11/9-12
- Sato: 11/9-12
- Hirata: 11/8-10